OpenStudy (starryknight):

please help me Find the 33rd term of the following sequence: 4, 9, 14 . . .

OpenStudy (anonymous):

so this is an AP with a=4 and d=5

OpenStudy (starryknight):

i got 164 for an an answer is that wrong or right

OpenStudy (anonymous):

An=a+(n-1)d A33=4+(33-1)5

OpenStudy (anonymous):

4+(32*5)=164

OpenStudy (anonymous):

you are absolutely right

OpenStudy (starryknight):

thank you ^^

OpenStudy (anonymous):

you are welcome